Economic model is characterised by the following data:
* Private consumption C = 800 + 0.9*DI
* Gross investment Ig = 400
* Government spending G = 500
* Sum of Taxes T = 300
* Disposable income DI = Y – T
Calculate:
* Equilibrium level of income Y (Y= 14300)
* Private consumption at macroeconomic equilibrium C = 13400

Develop equation of saving and calculate amount of saving at the point of macroeconomic equilibrium. S=. 600
Task 4. Economic model is characterised by the following data: Private consumption C-800+0.9*DI Gross investment Ig 400 Government spending G 500 Sum of Taxes T 300 Disposable income DI-Y T Calculate: a) Equilibrium level of income Y (Y 14300) b) c) Private consumption at macroeconomic equilibrium C 13400 Develop equation of saving and calculate amount of saving at the point of macroeconom equilibrium. S. 600
Answer #1

A.

Y = C + I + G

Y = 800+.9*DI + 400+ 500 = 800+.9*(Y-300) + 400+ 500

.1*Y = 1700 + .9Y - 270

.1Y = 1430

Y = 1430/.1

Y = $14300

B.

Private consumption = 800 + .9*DI

Private consumption = 800 + .9*(Y-T) = 800 + .9*(14300-300)

Private consumption = $13400

C.

Savings = income - private consumption - taxes = 14300-13400-300

Savings = $600